Adiusting Entry Practice Problem HINT: Before you make the adjusting entry, be sure to determine if they initially recorded that transactions under the regular approach or the alternative approach. Looking at the trial balance should help you make this determination. Welte and Moore Company are in the process of preparing their financial statements for the MONTH ended December 31, 2009. The Company has provided you with the following selected account balances form its UNADJUSTED trial balance: DEBIT CREDIT Cash s 20,000 18,000 125,000 $10,000 Insurance Expense Supplies Expense Service Revenue Accounts Payable Unearned Service Revenue 36,000 15,000 110,000 20,000 60,000 Additional Data: At the end of the year it was determined that $5,000 of supplies were still on hand. The employees have earned money for work done during December, but for which they will not be paid until January 2010. The amount was $3,000. The Company paid their fire insurance on December 1, 2009 for 9 months in advance. (a) (b) (c) (d) The machinery is expected to have a useful life of 10 years with an estimated scrap value (e) Customers paid the Company $60,000 in advance for services to be performed equally REQUIRED: ON THE NEXT PAGE PREPARE THE PROPER ADJUSTING ENTRIES of $5,000. over 3 months- December 2009, January (2010) and February (2010). REQUIRED AT 12-31-09. SHOW ALL COMPUTATIONS CLEARLY!!! USE THE FORMAT PROVIDED. The company adjusts its books monthly!!!
media%2Ffa2%2Ffa2bc657-8063-4c0a-9a60-44
Answer #1
ADJUSTING ENTRY Debit Credit COMPUTATION
a) Supplies 5000 No calculation is needed. The entire purchases of $15000 has
Supplies expense 5000 been expensed initially, without having a 'Supplies account'.  
Hence, the balance of stock of supplies in hand of $5000 needs  
to be brought into books and in the process, the supplies expense
has also got to be reduced to the extent of $5000.
b) Wages and salaries expense 3000 No computation is needed as the amount of unpaid salaries and
Wages and salaries payable 3000 wages is given.
c) Insurance prepaid 32000 The 9 months premium has been expensed on payment, instead of  
Insurance expense 32000 debiting to Prepaid insurance. Hence, 8 months premium is  
unexpired. This has to be shown as 'Prepaid insurance' and the
same amount has to reduced from insurance expenses.
Amount = 36000*8/9 = $32000
d) Depreciation expense 1000 Annual depreciation = (125000-5000)/10 = 12000
Accumulated depreciation 1000 Monthly depreciation = 12000/12 = $1000
e) Unearned service revenue 20000 Revenue for 1 month = 60000/3 = $20,000
Service revenue 20000